OpenStudy (anonymous):

If h(x) = −3x − 10, find h(−3) is the answer for this 1

OpenStudy (lgbasallote):

h(-3) = -3(-3) - 10 h(-3) = 9 - 10 check again ;)

OpenStudy (anonymous):

h(x) = −3x − 10 so h(-3) = -3*(-3) -10

OpenStudy (anonymous):

-3(-3)-10=-1

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

PUT x=-3 -3(-3)-10 9-10 -1